OpenStudy (anonymous):

In Arithmetwists (by Marcy Crook ) In the equations section... There is a problem i can't put my head around. It says make an equation with the digits 3,6, and 7. Each digit can only be used once. =21

OpenStudy (anonymous):

This is sooo confusing I am sorry.

OpenStudy (asnaseer):

if it doesn't say that ALL digits MUST be used, then you can say: 7 * 3 = 21

jimthompson5910 (jim_thompson5910):

or you can say 7*(6-3)

OpenStudy (asnaseer):

I believe jim is the winner!

OpenStudy (turingtest):

how did I not see that?

OpenStudy (anonymous):

No that equals 14

jimthompson5910 (jim_thompson5910):

7*(6-3) 7*3 21

OpenStudy (asnaseer):

7 * (6-3) = 7 * 3 = 7 + 7 + 7 = 21

OpenStudy (anonymous):

I am sorry I am olny in 5th grade... I don't know what you're talking about.

OpenStudy (asnaseer):

ok. do you agree that: 6 - 3 = 3

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

and do you agree that: 7 * 3 = 21

OpenStudy (anonymous):

I dont know what * means

jimthompson5910 (jim_thompson5910):

oh that's multiplication

OpenStudy (asnaseer):

\[7\times3=21\]

jimthompson5910 (jim_thompson5910):

7 * 3 is the same as 7 times 3

OpenStudy (anonymous):

Ohhhh. I feel stupid. Thank you

OpenStudy (asnaseer):

np - we have /moments/ like this :)

jimthompson5910 (jim_thompson5910):

no worries, if you've never seen it before then you're not stupid

OpenStudy (asnaseer):

*we ALL have...

OpenStudy (anonymous):

:) you are awesome

OpenStudy (asnaseer):

:) thx

OpenStudy (anonymous):

HOw about 24?

OpenStudy (asnaseer):

24 = 6 * (7-3)

OpenStudy (anonymous):

thx now i am done!!!

OpenStudy (asnaseer):

great!